OpenStudy (anonymous):

find an equation of the plane passing through the point(-2,3,4) and is perpendicular to the line passing through the points (4,-2,5) and (0,2,4)

OpenStudy (turingtest):

to define a plane we need a normal vector\[\vec n\cdot(\vec r-\vec r_o)=0\]and a point in the plane since we are given that the line passing through (4,-2,5) and (0,2,4) must be perpendicular to the plane we will use that to find our normal vector so first step, what is that vector?

OpenStudy (turingtest):

@bigboss still there?

OpenStudy (anonymous):

the dont provide a vector. i think u need to solve my previous questions first?

OpenStudy (anonymous):

1. find the plane that passes through the point(2,4,-3) and is parallel to the plane -2x+4y-5z+6=0

OpenStudy (anonymous):

2. find the line that passes through the point (2,5,3) and is perpendicular to the plane 2x-3y+4z+7 = 0

OpenStudy (turingtest):

you don't need any more info to solve this problem which one do you want to do first?

OpenStudy (anonymous):

can we start with 1, 2 and then the equation question?

OpenStudy (turingtest):

1. find the plane that passes through the point(2,4,-3) and is parallel to the plane -2x+4y-5z+6=0 ^this is what we are going to do first then.... did you see the hint I gave you about finding the normal vector when given the equation of a tangent plane on your other post?

OpenStudy (turingtest):

Two planes are parallel if their normal vectors are perpendicular, so we need the normal vector \(\vec n\) for the plane \(-2x+4y-5z+6=0\) if we are to construct a parallel plane

OpenStudy (turingtest):

*...if their normal vectors are parallel

OpenStudy (turingtest):

so what is the normal vector of \(-2x+4y-5z+6=0\)? hint: for plane equations of the form\[ax+by+cz=d\]the normal vector is given by\[\vec n=<a,b,c>\]

OpenStudy (anonymous):

ok im with you

OpenStudy (turingtest):

great, so then you know we need \(\vec n\) from \(-2x+4y-5z+6=0\) what is it?

OpenStudy (anonymous):

n = <-2,4,5> ?

OpenStudy (turingtest):

careful\[\vec n=<-2,4,-5>\]

OpenStudy (anonymous):

correct, sorry

OpenStudy (turingtest):

let \(\vec r=<x,y,z>\) be the position vector for any point in the plane let the position vector for our specific known point be denoted by\[\vec r_0=<2,4,-3>\]if you think about it, this means that \(\vec r-\vec r_0\) will be a vector inside our plane. Now, remembering that the dot product between two perpendicular vectors will be zero, we have that\[\vec n\cdot(\vec r-\vec r_0)=0\implies \vec n\cdot\vec r=\vec n\cdot\vec r_0\]do out the dot product above and you will get the equation of the plane

OpenStudy (anonymous):

can you do that for me? im a lil lost here

OpenStudy (turingtest):

I think you need to understand the derivation of the formula for a plane, which is well-described here: http://tutorial.math.lamar.edu/Classes/CalcIII/EqnsOfPlanes.aspx please do read that link when we are through here, but I will help you more with this one for now....

OpenStudy (anonymous):

thanks alot

OpenStudy (turingtest):

we agree that\[\vec n=<-2,4,-5>\]right? and do you agree that any other plane that has that same normal vector will be parallel to this one?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

I now am going to define two position vectors: (those are vectors that point from the origin to the point in question) \(\vec r=<x,y,z,>\) will be what I call the vector that points to any point \(P\) in the plane. We will have to leave the components as variables for now. The other position vector points to the specific point \(P_0\) we know the plane must pass through. Let's call it \(\vec r_0=<2,4,-3>\). (please look at the picture in the link so you know what I mean by these position vectors) -let me know when you have gotten all this....

OpenStudy (anonymous):

okay im listening

OpenStudy (turingtest):

so can you see from the picture in the link that \(\vec r-\vec r_0\) is in the plane?

OpenStudy (turingtest):

(it is red in the picture)

OpenStudy (anonymous):

yes i see it

OpenStudy (turingtest):

now since the normal vector \(\vec n\) is perpendicular to the plane, and because the dot product of two perpendicular vectors is zero, we know that\[\vec n\cdot(\vec r-\vec r_0)=0\]which means that\[\vec n\cdot\vec r=\vec n\cdot\vec r_0\]still with me? (fear not, we are almost done)

OpenStudy (anonymous):

yess with u

OpenStudy (turingtest):

so now just do those dot products above, and out pops the equation you are looking for you know what \(\vec n, \vec r,\) and \(\vec r_0\) are so we can compute\[\vec n\cdot\vec r=\vec n\cdot \vec r_0\]just plug in and do the dot products think you can do that? (it's just a plug-in basically)

OpenStudy (anonymous):

(-2+4+-5)*(2+4+-3) ?

OpenStudy (turingtest):

that almost looks like the right hand side of what we need, but why did you put + signs between all the components? you should write them with commas (-2,4,-5)*(2,4,-3)=? what is the result of this dot product?

OpenStudy (anonymous):

(-2,4,-5)*(2,4,-3)

OpenStudy (turingtest):

do you not know how do do dot products, i.e. scalar multiplication of vectors?

OpenStudy (turingtest):

how to do*

OpenStudy (turingtest):

if you do not know that dot products work like this\[<a,b,c>\cdot<x,y,z>=ax+by+cz\]then you missed something very important, and I do not see how you can do these problems...

OpenStudy (turingtest):

...but even if you have never worked with dot-products before, since I have given the formula for how to do them above you should now be able to tell me\[<-2,4,-5>\cdot<2,4,-3>=?\]

OpenStudy (anonymous):

(-2)(2)+(4)(4)+(-5)(-3)

OpenStudy (turingtest):

great, which equals...?

OpenStudy (anonymous):

27

OpenStudy (turingtest):

lovely so that is the RHS of our equation just to run through it again quickly, we found that\[\vec n\cdot\vec r=\vec n\cdot\vec r_0\]\[\vec n\cdot\vec r=<-2,4,-5>\cdot<2,4,-3>\]\[\vec n\cdot\vec r=27\]now what about the LHS ? what are those vectors and their dot-product?

OpenStudy (anonymous):

bro, i have to go now,thanks so much for ur help, can you please leave the answer for number 2 and 3? will look at it when im back

OpenStudy (turingtest):

nope, sorry

OpenStudy (anonymous):

LHS = 27 as well

OpenStudy (turingtest):

no it does not, I will show you this one, please do not leave

OpenStudy (anonymous):

ok plz show me

OpenStudy (turingtest):

\[\vec n\cdot\vec r=<-2,4,-5>\cdot<x,y,z>=-2x+4y-5z\]if the LHS was 27 we would have just gotten 27=27, which is definitely not the equation of a plane so in total we have...

OpenStudy (turingtest):

\[\vec n\cdot\vec r=\vec n\cdot\vec r_0\]\[<-2,4,-5>\cdot<x,y,z>=<-2,4,-5>\cdot<2,4,-3>\]\[-2x+4y-5y=27\]

OpenStudy (anonymous):

ok i get u nw

OpenStudy (anonymous):

so thats the final answer?

OpenStudy (turingtest):

...which is the equation of a plane, as opposed to 27=27 which is just a tautology so the steps are 1) find the normal vector to the given plane \(\vec n\) 2) define position vectors for a general point \(\vec r\), and a specific point in the plane \(\vec r_o\) 3)apply the formula \(\vec n\cdot\vec r=\vec n\cdot\vec r_0\)

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

can u help me wit 2 and 3 quick?

OpenStudy (turingtest):

and yes, the final answer is\[-2x+4y-5y=27\]not to be rough, but you really should be able to recognize that fact by the time we arrive at it more important is that you understand what we did here so I don't have to explain every step in the other problems, because I will \(not\) just give the answer so yeah, let's do 2 and 3, but please post a separate thread, this one is getting a bit long

OpenStudy (anonymous):

ok hold on, ill post a new thread